Hannah is mixing two types of Easter candy. The chocolates cost $5 a pound, and the caramels cost $4 a pound. If she creates 8 pounds worth $4.25 a pound, how much of the chocolate did she use?
Found 2 solutions by richwmiller, amalm06:
Answer by richwmiller(17219)

h= chocolates
a= caramel
5h+4a=8*4.25
a+h=8
a=6
h=2

Answer by amalm06(224)

Using the method of alligation, let CH=5, CR=4.
Then 4.25-4=0.25 and 5-4.25=0.75
So that CH/CR=0.25/0.75=(1/4)/(3/4)=1/3
Then the proportion of chocolates is 1/(1+3)=1/4
So that 8*(1/4)=2 chocolates (Answer)